OpenStudy (anonymous):

I have four identical oranges. How many ways are there for me to divide these oranges into at most three groups? (By definition, a group must have at least one orange.)

OpenStudy (ybarrap):

Group 1 - all groups of 1 orange. There are only 4 oranges, so 4 groups Group 2 - all groups of 2 oranges. 4 oranges arranged in groups of 2 is 4 choose 2 Group 3 - all groups of 3 oranges. 4 oranges arranged in groups of 3 is the same as 4 oranges arranged in groups of 1 Add all these 3 possibilities.

OpenStudy (anonymous):

Well, wouldn't group 1 not count? Because it says at most 3 groups?

OpenStudy (anonymous):

Also, I think it means to put them into groups simultaneously, and with no more than 3 groups

OpenStudy (ybarrap):

There are not many ways to make groups. Only group 4 is left.

OpenStudy (anonymous):

@satellite73 If you have time, can you help me on this one?

OpenStudy (anonymous):

dunno, too many glasses of wine but we can try it 1 group, 1 way two groups i would say 4 ways, but that might be wrong because it says you cannot tell them apart

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

three groups i would say \(4\times 3\times 2\times 3\) but i am confused about telling them apart

OpenStudy (anonymous):

Okay, I've confirmed that it rearrangements of the same way are still considered the same. i.e. 1,3 3,1 would count as 1 way, not two.

OpenStudy (anonymous):

theorem two tells us \(\dbinom{3+4-1}4=\dbinom64=\dfrac{6!}{2!4!}=15\)

OpenStudy (anonymous):

How did you get 3 and 4? I don't quite understand the second theorem.

OpenStudy (anonymous):

oh permutations are identical?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Which is why I'm thinking it is just- 1 Group- 4 balls all together, 1 way 2 Groups- 1,3 (3,1) 2,2 (2,2) 2 ways 3 Groups - 1,1,2 (2,1,2;2,1,1) 1 way.... Am I missing something?

OpenStudy (anonymous):

@wgary essentially we're counting the number of ways we can place 'bars' to separate the oranges: